QUESTION 1

A crate at rest resting on a floor:

Normal force direction will be:

along the angle of the ramp with respect to the horizontal

towards the center of earth

parallel to the surface of the ramp

perpendicular to the surface of the ramp

1 points   

QUESTION 2

A box is at rest on a wooden surface.

The frictional force direction in this case will be:

Opposite to the direction of the applied force

perpendicular to the surface

along the direction of the applied force

towards the center of the earth

1 points   

QUESTION 3

A crate being pushed and sliding up the ramp.

Normal force direction will be:

parallel to the ramp and upwards

parallel to the ramp and downwards

prependicular to the ramp and downwards

prependicular to the ramp and upwards

1 points   

QUESTION 4

A crate is being pushed and sliding upwards.

Frictional force direction:

prependicular to the ramp and downwards

parrallel to ramp pointing upwards

prependicular to the ramp and upwards

parrallel to ramp pointing downwards

1 points   

QUESTION 5

For every force there is a reaction force which is equal in magnitude. this law is called _____________

Law of Inertia

Newton's Third Law

Newton's Second Law

Newton's First Law

1 points   

QUESTION 6

Unit of weight in SI unit is called ___________

m/s

m/sec2

Newton

kilogram

1 points   

QUESTION 7

Unit of Force in SI unit is called ___________

Mass

kilogram

Inertia

Newton

1 points   

QUESTION 8

The vector sum of all the forces acting on the object is called __________

Normal force

Net force

Frictional force

Vector force

1 points   

QUESTION 9

A cubical box slides across a rough table. If the co-efficient of kinetic friction between the box and table is 0.3, what is the acceleration of the box?

(Hint: Use force equations and take the applied force = 0 for this case)

9 m/s2

12 m/s2

6 m/s2

3 m/s2
Answer #1

Q1.

Normal force will always be "perpendicular to the surface of the ramp"

Q2.

Friction force will be "Opposite to the direction of the applied force"

Q3.

Again normal force will be "Perpendicular to the surface of ramp"

Q4.

Since box is pushing upwards, So friction force will be in opposite direction, which is "parrallel to ramp pointing downwards".

Q5.

It's Newton's third law

Q6.

SI unit of weight = Newton

Q7.

SI unit of Force = Newton

Q8.

The vector sum of all the forces acting on the object is called "Net Force".

Q9.

Fnet = Ff, since there is No other force

Ff = uk*N

N = m*g

Fnet = m*a

m*a = uk*m*g

a = uk*g

a = 0.3*10 = 3 m/sec^2

Correct option is D.